A men's store bought 500 suits, some at $125 each and the rest at $200 each. If the total cost of the suits was $77,500, how many suits was purchased at $200?


1
Expert's answer
2021-04-29T16:04:19-0400

Let the number of suits bought at $125 be x


Let the number of suits bought at $200 be y




Fom the question we derived two equations:


125(x) + 200(y) = 77500........... Equation 1

x + y = 500.......... Equation 2



From equation 2 [ x + y = 500]


We can say x = 500 - y




Since x = 500 - y




Input 500 - y for the value of x in equation 1




∴ 125(500 - y) + 200(y) = 77500

62500 - 125y + 200y =77500

62500 + 75y = 77500

75y = 77500 - 62500

75y = 15000



Divide both sides by 75




∴ y = 200

y represents the number of suits bought at $200




∴ the number of suits bought at $200 is 200